# EnvPoser: Environment-aware Realistic Human Motion Estimation from Sparse Observations with Uncertainty Modeling

### v1

@@ Abstract
Estimating full-body motion using the tracking signals of head and hands from VR devices holds great potential for various applications. However, the sparsity and unique distribution of observations present a significant challenge, resulting in an ill-posed problem with multiple feasible solutions (i.e., hypotheses). This amplifies uncertainty and ambiguity in full-body motion estimation, especially for the lower-body joints. Therefore, we propose a new method, EnvPoser, that employs a two-stage framework to perform full-body motion estimation using sparse tracking signals and pre-scanned environment from VR devices. EnvPoser models the multi-hypothesis nature of human motion through an uncertainty-aware estimation module in the first stage. In the second stage, we refine these multi-hypothesis estimates by integrating semantic and geometric environmental constraints, ensuring that the final motion estimation aligns realistically with both the environmental context and physical interactions. Qualitative and quantitative experiments on two public datasets demonstrate that our method achieves state-of-the-art performance, highlighting significant improvements in human motion estimation within motion-environment interaction scenarios.  [^1] Code will be released upon paper acceptance.
@@ 1 Introduction
Our key insight is that joint uncertainty estimation can explicitly model the multi-hypothesis nature of full-body motion estimation, and by incorporating environmental information, we can reduce estimation uncertainty, guiding the motion estimation to converge toward the most plausible outcome that aligns with both the environment and sparse input observations. Therefore, we design a two-stage deep learning framework to handle the one-to-many mapping ambiguity challenge. In the first stage, we explicitly model the multi-hypothesis nature of human motion by the joint uncertainty estimation and its sampling strategy. Then, by integrating environmental information, we guide the uncertainty-based multi-hypothesis motion estimations to converge on a full-body motion that best aligns with both the environment and sparse observations. To address the challenge of effectively utilizing environmental information, we refine motion estimation using semantic and geometric constraints for non-contact and contact scenarios, respectively. Semantic information provides contextual cues for realistic estimation in non-contact motions, while geometric constraints prevent collisions in contact scenarios, correcting unrealistic motions.
@@ 2.1 Motion estimation with sparse observations
While methods utilizing six IMUs have shown impressive results, HMD-based techniques leveraging sparse three-point observations show significant potential for practical applications. For instance, AvatarPoser [14] employs a transformer architecture to achieve accurate full-body motion. While Zheng et al. [54] developed a two-stage framework to capture joint dependencies, improving motion accuracy and fluidity. Some studies [5, 1, 8] address the ill-posed problem of sparse-to-dense mapping using generative models. Dittadi et al. [5] and Aliakbarian et al. [1] proposed using Variational Autoencoders (VAE) and normalizing flows, respectively, to estimate full-body poses from three tracking points. Additionally, the diffusion model [6, 8, 34] has also shown notable performance in motion estimation with sparse data.
Despite these advances, most methods rely solely on sparse tracking signals, leading to high uncertainty and multiple potential poses for the same input. Therefore, this paper introduces joint uncertainty estimation and constrains physically implausible motions using environmental information.
@@ 3.1 Problem Statement
Our goal is to estimate the full-body motion sequence $\boldsymbol{\theta}=\{\theta_{t}\}^{T}_{t=1}\in\mathcal{R}^{T\times 132}$ using the sparse tracking signals $\boldsymbol{X}=\{{\boldsymbol{x}}_{t}\}^{T}_{t=1}\in\mathcal{R}^{T\times N_{c}}$ from the HMD and hand controllers, along with a pre-scanned 3D environment point cloud $\boldsymbol{S}$, where $T$ denotes the length of input, $N_{c}$ = 36 means the channel of input. Each sparse tracking signal at time step $t$ consists of the head and hand positions $\boldsymbol{p}_{t}$, rotations $\boldsymbol{R}_{t}$, and linear velocities $\boldsymbol{v}_{t}$, represented as $\boldsymbol{x}_{t}=[\boldsymbol{p}_{t},\boldsymbol{R}_{t},\boldsymbol{v}_{t}]$. Moreover, we utilize a cropped 3D environment point cloud $\boldsymbol{V_{S}}\in\mathcal{R}^{N_{S}\times 3}$, where $N_{S}$, the number of sampled points, is set to 1000. Adjusted based on human translation, this point cloud refines pose estimation by incorporating environmental context, leveraging both contact and non-contact interactions between the body and the environment to reduce the uncertainty inherent in sparse motion observations.

### v2

@@ Abstract
Estimating full-body motion using the tracking signals of head and hands from VR devices holds great potential for various applications. However, the sparsity and unique distribution of observations present a significant challenge, resulting in an ill-posed problem with multiple feasible solutions (i.e., hypotheses). This amplifies uncertainty and ambiguity in full-body motion estimation, especially for the lower-body joints. Therefore, we propose a new method, EnvPoser, that employs a two-stage framework to perform full-body motion estimation using sparse tracking signals and pre-scanned environment from VR devices. EnvPoser models the multi-hypothesis nature of human motion through an uncertainty-aware estimation module in the first stage. In the second stage, we refine these multi-hypothesis estimates by integrating semantic and geometric environmental constraints, ensuring that the final motion estimation aligns realistically with both the environmental context and physical interactions. Qualitative and quantitative experiments on two public datasets demonstrate that our method achieves state-of-the-art performance, highlighting significant improvements in human motion estimation within motion-environment interaction scenarios. Project page: https://xspc.github.io/EnvPoser/.
@@ 1 Introduction
Our key insight is that joint uncertainty estimation can explicitly model the multi-hypothesis nature of full-body motion estimation, and by incorporating environmental information, we can reduce estimation uncertainty, guiding the motion estimation to converge toward the most plausible outcome that aligns with both the environment and sparse input observations. Therefore, we design a two-stage deep learning framework to handle the one-to-many mapping ambiguity challenge. We first explicitly model the multi-hypothesis human motion by the joint uncertainty estimation and its sampling strategy. Then, by integrating environmental information, we guide the uncertainty-based multi-hypothesis motion estimations to converge on a full-body motion that best aligns with both the environment and sparse observations. To address the challenge of effectively utilizing environmental information, we introduce semantic constraints for non-contact motions, providing contextual cues for realistic estimation, and geometric constraints for contact scenarios, ensuring consistency and preventing collisions.
@@ 2.1 Motion estimation with sparse observations
While methods utilizing six IMUs have shown impressive results, HMD-based techniques leveraging sparse three-point observations show significant potential for practical applications. For instance, AvatarPoser [14] employs a transformer architecture to achieve accurate full-body motion. While Zheng et al. [55] developed a two-stage framework to capture joint dependencies, improving motion accuracy and fluidity. Some studies [5, 1, 8] address the ill-posed problem of sparse-to-dense mapping using generative models. Dittadi et al. [5] and Aliakbarian et al. [1] proposed using Variational Autoencoders (VAE) and normalizing flows, respectively, to estimate full-body poses from three tracking points. Additionally, the diffusion model [6, 8, 35] has also shown notable performance in motion estimation with sparse data. Despite these advances, most methods rely solely on sparse tracking signals, leading to high uncertainty and multiple potential poses for the same input. Therefore, this paper introduces joint uncertainty estimation and constrains physically implausible motions using environmental information.
@@ 3.1 Problem Statement
Our goal is to estimate the full-body motion sequence $\boldsymbol{\theta}=\{\theta_{t}\}^{T}_{t=1}\in\mathcal{R}^{T\times 132}$ (the first 22 joints of SMPL [26] model with the 6D representation of rotations) using the sparse tracking signals $\boldsymbol{X}=\{{\boldsymbol{x}}_{t}\}^{T}_{t=1}\in\mathcal{R}^{T\times N_{c}}$ from the HMD and hand controllers, along with a pre-scanned 3D environment point cloud $\boldsymbol{S}$, where $T$ denotes the length of input, $N_{c}$ = 36 means the channel of input. Each sparse tracking signal at time step $t$ consists of the head and hand positions $\boldsymbol{p}_{t}$, rotations $\boldsymbol{R}_{t}$, and linear velocities $\boldsymbol{v}_{t}$, represented as $\boldsymbol{x}_{t}=[\boldsymbol{p}_{t},\boldsymbol{R}_{t},\boldsymbol{v}_{t}]$. Moreover, we utilize a cropped 3D environment point cloud $\boldsymbol{V_{S}}\in\mathcal{R}^{N_{S}\times 3}$, where $N_{S}$, the number of sampled points, is set to 1000. With the human translation, this point cloud refines pose estimation by incorporating environmental context, leveraging both contact and non-contact interactions between the body and the environment to reduce the uncertainty inherent in sparse observations.